The surface area of a cube is 24 cm2. what is the new surface area of the cube if the length of each edge of the cube is tripled.
Answer by colliefan(239)

Let's start by finding the length of the edge of the initial cube. A cube has 6 different surfaces that are all squares that have the same length and widths and so have the same areas. The surface area of the cube must be 6 * the area of one square because there are six surfaces. As an equation, if S is surface area and a is area of one square: s=6a.
We know that s=24 cm^2 for the initial cube. Putting that in to the equation,
24=6a
24/6=6a/6
4=a
Now we know the area of one square to be 4 cm^2. If the area of the square is 4 cm^2, the length of one side must be 2 since the area of a square is the length squared. a=l^2
4=l^2
2=l
______
If the initial length is 2, tripling it gives a square with length 6. The area of a square with length 6 is 6^2 or 36. The surface area of the cube made up of such squares is 6*36 or 216.